# Telegrapher’s Generative Model via Kac Flows

### v5

@@ 2 Preliminaries: Curves in Wasserstein Spaces
$W_{2}(\mu_{s},\mu_{t})\leq\int_{s}^{t}m(r)\,\,\mathrm{d}r\quad\text{for all }s\leq t,\,s,t\in I.$ (2.1)
Here $L_{2}(I)$ denotes the space of (equivalence classes of) square integrable real-valued function on an interval $I$. It is well known that a narrowly [^1] In stochastic analysis, narrow convergence is also known as weak convergence of measures. continuous curve $\mu_{t}:I\to\mathcal{P}_{2}(\mathbb{R}^{d})$ belongs to $AC^{2}(I;\mathcal{P}_{2}(\mathbb{R}^{d}))$, iff there exists a Borel measurable vector field $v:I\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $(\mu_{t},v_{t})$ satisfies the continuity equation
@@ 3 Motivation
Learning absolutely continuous curves induced by couplings of the latent and target measure is a task addressed in flow matching, see [27], and for an overview, [47]. Let us consider the latent $\mu_{0}=\mathcal{N}(0,I_{d})\,\mathrm{d}x$ and the target distribution $\mu_{1}=\delta_{0}$, and their independent coupling $\alpha:=\mu_{0}\times\mu_{1}$. Then the curve $\mu_{t}:=e_{t,\sharp}\alpha$ with $e_{t}=(1-t)x+ty$ admits the density
$p_{t}(x)$ $=\left(2\pi(1-t)^{2}\right)^{-\frac{d}{2}}{\rm e}^{-\frac{\|x\|^{2}}{2(1-t)^{2}}},\quad t\in I=[0,1)$ (3.1)
and with the so-called velocity field induced by $\alpha$,
@@
see [47, Example 4.12], it fulfills the continuity equation (2.2). Unfortunately, (2.4) is not fulfilled due to the singularity at $t=1$. Nevertheless, the trajectory $\varphi_{t}$ of the ODE (2.5) is given by
@@
Diffusion models or score-based models are based on the Fokker-Planck equation of the Wiener process or Ornstein-Uhlenbeck process, which is the diffusion equation (with drift), see e.g. [41]. Let us consider the probability flow $p_{t}$ solving the standard diffusion equation
@@
$\tilde{p}_{t}(x)=\big(2\pi(1-t)\big)^{-\frac{d}{2}}{\rm e}^{-\frac{\|x\|^{2}}{2(1-t)}},\quad t\in I=[0,1),\quad\tilde{p}_{0}=\mathcal{N}(0,I_{d}),$ (3.5)
@@
$\partial_{t}p_{t}=\nabla\cdot(p_{t}v_{t})=\left(-\langle\nabla f_{t},v_{t}\rangle+\nabla\cdot v_{t}\right)p_{t},$ (3.9)
@@
Sampling of unnormalized target Boltzmann densities $p_{1}$ often uses the geometric density interpolation $p_{t}=p_{0}^{1-t}p_{1}^{t}$ which corresponds to the linear interpolation $f_{t}=(1-t)f_{0}+tf_{1}$. Then (3.10) reads as
@@
Note that replacing the linear interpolation by a more flexible variant with an additional learnable quantity, the mode collapse was numerically avoided in [máté2023learning]. For an approach to (3.10) from the control theory point of view that leverages the explosion effect, we refer to [11, 4, 43].
@@ 4 Telegrapher’s Equation and Kac Model in 1D
In [28], it was contemplated to consider other physics-inspired models next to above diffusion (parabolic) and Poisson (elliptic) equations for generative modeling. Here, we further pursue this idea and focus on a process related to the telegrapher’s equation, also known as damped wave equation. This is a linear hyperbolic PDE of the form
@@
Here, $a>0$ is the damping coefficient, and $c>0$ describes the velocity of the wave front. Historically, the equation (4) originates from the modeling of voltage and current in electrical transmission lines (telegraphs). It is known that if the damping $a$ and velocity $c$ go suitably to infinity, see Remark 5.5, then the diffusion model is retrieved. Hence, diffusion can be seen as "an infinitely $a$-damped wave with infinite propagation speed cc". The idea of particles traveling with infinite speed has already encountered resistance in the physics community [7, 46, 10, 31, 44], since it violates Einstein’s laws of relativity. But also in the field of generative modeling, this can be seen as a reason for the explosion of velocity fields of diffusion models. Hence, if we restrict ourselves to above model (4), where waves travel only with finite speed $c$, we expect the velocity field $v_{t}$ to behave well. Indeed, we will show that the telegrapher’s model does not admit a velocity explosion, but shares the same long-time asymptotics as diffusion.
@@ 4.1 Kac Flow in 1D
Suppose that a particle starts at $x_{0}=0$ and always moves with velocity $c>0$ in either positive or negative direction. The initial direction is determined by a random symmetric coin-flip. Each step is of duration $\Delta t$ and therefore we have $\Delta x=c\Delta t$. A particle continues to move in the same direction with probability $1-a\Delta t$, and reverses with the probability $a\Delta t$, where $a>0$ is a fixed parameter. Intuitively, for small $\Delta t$ the particle tends to move in the same direction (resembling some form of inertia in the movement), until it collides with some obstacle and reverses its direction.
@@
We refer to $\tau_{t}$ as random time. Finally, for a (symmetrically) random initial direction, the displacement at time $t$ is determined by the Kac process starting in 0,
@@
where ${\rm B}_{\frac{1}{2}}\sim B(1,\frac{1}{2})$ is a Bernoulli random variable taking the values $\pm 1$. Naturally, for a random variable $X_{0}$ we call
@@ 4.2 Relation of Kac’ Flow to 1D Telegrapher’s Equation
By the next lemma, there is an analytic representation of the distributional solution of the telegrapher’s equation (4). We refer to [34] and [31] for the proof via distributions. For a characterization of measures as (tempered) distributions, see also [38, Chapter 4.4].
@@
Let $f_{0}=\delta_{0}$ be the Dirac measure on $\mathbb{R}$. Then $\mu_{t}$ associated to $u(t,\cdot)$ in (4.13) solves the continuity equation
@@ 5 Multi-dimensional Generalizations of the Telegrapher’s
The connection between the telegrapher’s equation and the Kac process can be generalized to higher dimensions via Kac’s insertion method which we briefly describe in Subsection 5.1. In contrast to one dimension, initial probability densities do not remain probability densities in the PDE formulation. Yet, we state a convergence result towards diffusion, which we will use subsequently. To circumvent the lack of conservation of mass, we propose two types of componentwise extensions to higher dimensions in Subsection 5.2, and we prove regularity and convergence results of our componentwise processes.
@@ 5.2 Multi-Dimensional Kac Process
    $\|v_{t}\|_{L_{2}(\mu_{t})}^{2}\leq dc^{2}\quad\text{ for a.e. }t\in[0,\infty),$ (5.13)
@@
ii) By definition (2.1) and Part i), we see immediately that $(\mu_{t})_{t\geq 0}$ is an absolutely continuous curve. Then it follows from [3, Theorem 8.3.1] that there is a (optimal) velocity field in the continuity equation fulfilling (5.13).
@@ 7 Numerical Experiments
In this section, we use the flow matching framework to train a neural network to approximate the velocity field from Corollary 6.3 by minimizing the CFM loss given in (6.13). To this end, we compute the conditional velocity $v_{t}(x|x_{0})$ via its analytic formula (6.9). In order to sample from $\mathbf{X}_{t}\sim u_{t}(\cdot|x_{0})$, where $x_{0}$ is fixed, we only need to sample i.i.d. components $X_{t}^{i}$ of the Kac process starting in 0. This can be done efficiently, since we have access to the analytic form (4.13) of the one-dimensional Kac distribution starting in 0 (use e.g. inverse transform sampling). Alternatively, if we wish to sample fixed realizations of a Kac process for varying times $t$, we can efficiently simulate the random walk itself as explained in Appendix A.3.
@@ 7.2 Image Generation
where $v_{K}(t,x\mid 0)$ is the conditional Kac velocity field from (6.4). Note that in the last equality, the independence of $X_{0}$ and $K(t)$ is used to drop the condition $X_{0}=x_{0}$. Now, by the same arguments as in Corollary 6.3, one recovers the corresponding vector field for the mean‐reverting Kac process.
@@ A.1 Lipschitz property of the Mean-reverting Kac process
Let us consider the $d$-dimensional versions $\mathbf{X}_{0}\coloneqq(X_{0}^{1},\dots,X_{0}^{d})$ and $\mathbf{M}_{t}\coloneqq(M_{t}^{1},\dots,M_{t}^{d})$ of the data and the mean-reverting Kac process (7.1), where only for notational simplicity, we assume the same scheduling functions, damping and speed parameters for each component. Similar to Proposition 5.6, we have the following Lipschitz property.
@@
Let $P_{\mathbf{X}_{0}}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, and let $f,g:[0,1]\to\mathbb{R}$ be Lipschitz continuous with constants $C_{f},C_{g}>0$, respectively. Then, for the probability distribution flow $(\mu_{t})_{t\geq 0}\coloneqq(P_{\mathbf{M}_{t}})_{t\geq 0}$ of the $d$-dimensional mean-reverting Kac process $\mathbf{M}_{t}$, the following holds true:
@@
    $(\mu_{t})_{t\geq 0}$ is Lipschitz continuous in the Wasserstein space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ with Lipschitz constant $\sqrt{2C_{f}^{2}\,\|\mathbf{X}_{0}\|_{L_{2}(\mathbb{P})}^{2}+2C_{g}^{2}\,dc^{2}}$, and for $\mathbb{P}$-a.e. $\omega$, $\mathbf{X}_{t}(\omega)$ is Lipschitz continuous in $\mathbb{R}^{d}$ with the constant $\sqrt{2C_{f}^{2}\,\|\mathbf{X}_{0}(\omega)\|^{2}+2C_{g}^{2}\,dc^{2}}$.
@@
    $\|v_{t}\|_{L_{2}(\mu_{t})}^{2}\leq 2C_{f}^{2}\,\|\mathbf{X}_{0}\|_{L_{2}(\mathbb{P})}^{2}+2C_{g}^{2}\,dc^{2}\quad\text{ for a.e. }t\in[0,\infty),$ (A.1)
@@
i) Fix $0\leq s<t$. Analogously as in (5.2), the construction of $\mathbf{M}_{t}$ via (7.1) yields for $\mathbb{P}$-a.e. $\omega$ that
@@
Hence, $(\mu_{t})_{t\geq 0}$ is Lipschitz continuous in $\mathcal{P}_{2}(\mathbb{R}^{d})$ with constant $(2C_{f}^{2}\,\|\mathbf{X}_{0}\|_{L_{2}(\mathbb{P})}^{2}+2C_{g}^{2}\,dc^{2})^{\frac{1}{2}}$.
ii) By definition (2.1) and Part i), we see immediately that $(\mu_{t})_{t\geq 0}$ is an absolutely continuous curve. Then it follows from [3, Theorem 8.3.1] that there is a (optimal) velocity field in the continuity equation fulfilling (A.1). ∎

### v6

@@ 2 Preliminaries: Curves in Wasserstein Spaces
$W_{2}(\mu_{s},\mu_{t})\leq\int_{s}^{t}m(r)\,\,\mathrm{d}r\quad\text{for all }s\leq t,~s,t\in I.$ (2.1)
Here $L_{2}(I)$ denotes the space of (equivalence classes of) square-integrable real-valued functions on $I$. It is well known that a narrowly [^1] In stochastic analysis, narrow convergence is also known as weak convergence of measures. continuous curve $\mu_{t}:I\to\mathcal{P}_{2}(\mathbb{R}^{d})$ belongs to $AC^{2}(I;\mathcal{P}_{2}(\mathbb{R}^{d}))$, iff there exists a Borel measurable vector field $v:I\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $(\mu_{t},v_{t})$ satisfies the continuity equation
@@ 3 Motivation
Learning absolutely continuous curves induced by couplings of the latent and target measure is a task addressed in flow matching, see [27], and for an overview, [47]. Let us consider the latent $\mu_{0}=\mathcal{N}(0,I_{d})$ and the target distribution $\mu_{1}=\delta_{0}$, and their independent coupling $\alpha:=\mu_{0}\times\mu_{1}$. Then the curve $\mu_{t}:=e_{t,\sharp}\alpha$ with $e_{t}=(1-t)x+ty$ admits the density
$p_{t}(x)$ $=\left(2\pi(1-t)^{2}\right)^{-\frac{d}{2}}{\rm e}^{-\frac{\|x\|^{2}}{2(1-t)^{2}}},\quad t\in[0,1),\quad{p}_{0}=\mathcal{N}(0,I_{d}),$ (3.1)
and with the velocity field induced by $\alpha$, see [47, Example 4.12],
@@
it fulfills the continuity equation (2.2). Unfortunately, (2.4) is not fulfilled due to the singularity at $t=1$. Nevertheless, the trajectory $\varphi_{t}$ of the ODE (2.5) is given by
@@
Diffusion models or score-based models are based on the Fokker-Planck equation of the Wiener process or Ornstein-Uhlenbeck process, i.e. the diffusion equation (with drift), see e.g. [41]. Let us consider the probability flow $p_{t}$ solving the standard diffusion equation
@@
$\tilde{p}_{t}(x)=\big(2\pi(1-t)\big)^{-\frac{d}{2}}{\rm e}^{-\frac{\|x\|^{2}}{2(1-t)}},\quad t\in[0,1),\quad\tilde{p}_{0}=\mathcal{N}(0,I_{d}),$ (3.5)
@@
$\partial_{t}p_{t}=-\nabla\cdot(p_{t}v_{t})=-\left(-\langle\nabla f_{t},v_{t}\rangle+\nabla\cdot v_{t}\right)p_{t},$ (3.9)
@@
Sampling from unnormalized target Boltzmann densities $p_{1}$ often uses the geometric density interpolation $p_{t}=p_{0}^{1-t}p_{1}^{t}$ which corresponds to the linear interpolation $f_{t}=(1-t)f_{0}+tf_{1}$. Then (3.10) reads as
@@
Note that by replacing the linear interpolation with a more flexible variant including an additional learnable term, the mode collapse was numerically avoided in [máté2023learning]. For a control-theoretic perspective on (3.10) that leverages the explosion effect, we refer to [11, 4, 43].
@@ 4 Telegrapher’s Equation and Kac Model in 1D
In [28], it was contemplated to consider other physics-inspired models next to above diffusion (parabolic) and Poisson (elliptic) equations for generative modeling. Here, we further pursue this idea and focus on a process related to the telegrapher’s equation [^2] Historically, the equation (4) originates from the modeling of voltage and current in electrical transmission lines (telegraphs)., also known as damped wave equation. This is a linear hyperbolic PDE of the form
@@
Here, $a>0$ is the damping coefficient, and $c>0$ describes the velocity of the wave front. It is known that if the damping $a$ and velocity $c$ go suitably to infinity, see Remark 5.5, then the diffusion model is retrieved. Hence, diffusion can be seen as "an infinitely $a$-damped wave with infinite propagation speed cc". The idea of particles traveling with infinite speed has already encountered resistance in the physics community [7, 46, 10, 31, 44], since it violates Einstein’s laws of relativity. But also in the field of generative modeling, this can be seen as a reason for the explosion of velocity fields of diffusion models. Hence, if we restrict ourselves to above model (4), where waves travel only with finite speed $c$, we expect the velocity field $v_{t}$ to behave well. Indeed, we will show that the telegrapher’s model does not admit a velocity explosion, but shares the same long-time asymptotics as diffusion.
@@ 4.1 Kac Flow in 1D
Suppose that a particle starts at $x_{0}=0$ and always moves with velocity $c>0$ in either positive or negative direction. The initial direction is determined by a random symmetric coin-flip. Each step is of duration $\Delta t$ and therefore we have $\Delta x=c\Delta t$. The particle continues to move in the same direction with probability $1-a\Delta t$, and reverses with the probability $a\Delta t$, where $a>0$ is a fixed parameter. Intuitively, for small $\Delta t$ the particle tends to move in the same direction (resembling some form of inertia in the movement), until it collides with some obstacle and reverses its direction.
@@
We refer to $\tau_{t}$ as random time. Finally, for a symmetric random initial direction, the displacement at time $t$ is determined by the Kac process starting in 0,
@@
where ${\rm B}_{\frac{1}{2}}\sim{\rm Ber}(\frac{1}{2})$ is a Bernoulli random variable taking the values $\pm 1$. Naturally, for a random variable $X_{0}$ we call
@@ 4.2 Relation of Kac’ Flow to 1D Telegrapher’s Equation
By the next lemma, there is an analytic representation of the distributional solution of the telegrapher’s equation (4). We refer to [31] and to [34] for the proof via distributions. For a characterization of measures as (tempered) distributions, see also [38, Chapter 4.4].
@@
Let $f_{0}=\delta_{0}$ be the Dirac measure at 0 in $\mathbb{R}$. Then $\mu_{t}$ associated to $u(t,\cdot)$ in (4.13) solves the continuity equation
@@ 5 Multi-dimensional Generalizations of the Telegrapher’s
The connection between the telegrapher’s equation and the Kac process can be generalized to higher dimensions via Kac’s insertion method, which we briefly describe in Subsection 5.1. In contrast to the one-dimensional case, initial probability densities do not remain probability densities in the PDE formulation. Yet, we state a convergence result towards diffusion, which we will use subsequently. To circumvent the lack of mass conservation, we propose two types of componentwise extensions in Subsection 5.2, and we prove regularity and convergence results for our componentwise processes.
@@ 5.2 Multi-Dimensional Kac Process
    $\|v_{t}\|_{L_{2}(\mu_{t})}^{2}\leq dc^{2}\quad\text{ for a.e. }t\in[0,\infty).$ (5.13)
@@
ii) By definition (2.1) and Part i), we see immediately that $(\mu_{t})_{t\geq 0}$ is an absolutely continuous curve. It then follows from [3, Theorem 8.3.1] that there exists an (optimal) velocity field in the continuity equation fulfilling (5.13).
@@ 7 Numerical Experiments
In this section, we use the flow matching framework to train a neural network to approximate the velocity field from Corollary 6.3 by minimizing the CFM loss given in (6.13). To this end, we compute the conditional velocity $v_{t}(x|x_{0})$ via its analytic formula (6.9). In order to sample from $\mathbf{X}_{t}\sim u_{t}(\cdot|x_{0})$, where $x_{0}$ is fixed, we only need to sample i.i.d. components $K_{t}^{i}$ of the Kac process starting in 0. This can be done efficiently, since we have access to the analytic form (4.13) of the one-dimensional Kac distribution starting in 0 (use e.g. inverse transform sampling). Alternatively, if we wish to sample fixed realizations of a Kac process for varying times $t$, we can efficiently simulate the random walk itself as explained in Appendix A.3.
@@ 7.2 Image Generation
where $v_{K}(t,x\mid 0)$ is the conditional Kac velocity field from (6.4). Note that in the last equality, the independence of $X_{0}$ and $K_{t}$ is used to drop the condition $X_{0}=x_{0}$. Now, by the same arguments as in Corollary 6.3, one recovers the corresponding vector field for the mean‐reverting Kac process.
@@ A.1 Lipschitz property of the Mean-reverting Kac process
Let us consider the $d$-dimensional versions $\mathbf{X}_{0}\coloneqq(X_{0}^{1},\dots,X_{0}^{d})$ and $\mathbf{M}_{t}\coloneqq(M_{t}^{1},\dots,M_{t}^{d})$ of the data and the mean-reverting Kac process (7.1). Only for notational simplicity, we assume the same scheduling functions, damping and speed parameters for each component, meaning that $\mathbf{M}_{t}=f(t)\mathbf{X}_{0}+\mathbf{K}_{g(t)}$. Similar to Proposition 5.6, we have the following Lipschitz property.
@@
Let $P_{\mathbf{X}_{0}}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, and let $f,g:[0,1]\to\mathbb{R}$ be Lipschitz continuous with constants $C_{f},C_{g}>0$, respectively. Then, for the probability distribution flow $(\mu_{t})_{t\in[0,1]}\coloneqq(P_{\mathbf{M}_{t}})_{t\in[0,1]}$ of the $d$-dimensional mean-reverting Kac process $\mathbf{M}_{t}$, the following holds true:
@@
    $(\mu_{t})_{t\in[0,1]}$ is Lipschitz continuous in the Wasserstein space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ with Lipschitz constant $\sqrt{2C_{f}^{2}\,\|\mathbf{X}_{0}\|_{L_{2}(\mathbb{P})}^{2}+2C_{g}^{2}\,dc^{2}}$, and for $\mathbb{P}$-a.e. $\omega$, $\mathbf{X}_{t}(\omega)$ is Lipschitz continuous in $\mathbb{R}^{d}$ with the constant $\sqrt{2C_{f}^{2}\,\|\mathbf{X}_{0}(\omega)\|^{2}+2C_{g}^{2}\,dc^{2}}$.
@@
    $(\mu_{t})_{t\in[0,1]}$ is an absolutely continuous curve in $AC^{2}([0,1];\mathcal{P}_{2}(\mathbb{R}^{d}))$, and admits a velocity field fulfilling
    $\|v_{t}\|_{L_{2}(\mu_{t})}^{2}\leq 2C_{f}^{2}\,\|\mathbf{X}_{0}\|_{L_{2}(\mathbb{P})}^{2}+2C_{g}^{2}\,dc^{2}\quad\text{ for a.e. }t\in[0,1].$ (A.1)
@@
i) Fix $0\leq s<t\leq 1$. Applying the estimate (5.2) and Young’s inequality, we obtain for $\mathbb{P}$-a.e. $\omega$ that
@@
Hence, $(\mu_{t})_{t\in[0,1]}$ is Lipschitz continuous in $\mathcal{P}_{2}(\mathbb{R}^{d})$ with constant $(2C_{f}^{2}\,\|\mathbf{X}_{0}\|_{L_{2}(\mathbb{P})}^{2}+2C_{g}^{2}\,dc^{2})^{\frac{1}{2}}$.
ii) By definition (2.1) and Part i), we see immediately that $(\mu_{t})_{t\in[0,1]}$ is an absolutely continuous curve. It then follows from [3, Theorem 8.3.1] that there exists an (optimal) velocity field in the continuity equation fulfilling (A.1). ∎